A motorist travelled at 90km/hour applied it's brake came to rest with uniform retardation in 20seconds, calculate the retardation in meter per seconds square



1
Expert's answer
2022-01-11T04:13:05-0500

Given:

"v_i=90\:\rm km/h=25\: m/s"

"v_f=0\:\rm km/h=0\: m/s"

"\Delta t=20\:\rm s"


The retardation

"a=\frac{\Delta v}{\Delta t}=\frac{0-25}{20}=-1.25\:\rm m/s^2"